# Science Time Series: Deep Learning in Hydrology

### v1

@@ III-B Model Setup
TABLE IV: Model Activation Setup
@@
III-C Model Training and Evaluation
Inputs to the model can be classified into known inputs and observed inputs [5]. Known inputs refer to features that are known in both the past and future. In this hydrology study, static known inputs are the exogenous features such as climatic signatures, hydrologic signatures, and catchment topography. Additionally, time series known inputs exist in the form of seasonal patterns hidden in all time series features. Observed inputs are features known only for the past time periods but unknown in the future. In hydrology, those features include precipitation, temperature, and streamflow. Model output, known as targets, are the time-dependent predicted properties. In this study, predicted targets are precipitation, temperature, and streamflow.
@@ III-C Model Training and Evaluation
Spatial and Temporal Encodings. Like many fields of science, hydrology time series data exhibit strong seasonal patterns. It can reasonably be assumed that, at a specific gauge location, the precipitation and streamflow in October of one year will be similar to those recorded in October of the previous year. Furthermore, research has identified the approximate water residence times in various reservoirs, such as rivers, lakes, soil, and the atmosphere [46]. To effectively capture these known dependencies within the time series properties, spatial and temporal encodings are incorporated into the model during training.
@@
NNSE. Model fit is quantified using normalized Nash-Sutcliffe Efficiency (NNSE) scores [47], calculated using Equation 11, where $T$ represents the total number of days, $Q_{m}^{t}$ represents the modeled discharge on day $t$, $Q_{o}^{t}$ represents the observed discharge on day $t$, and $\overline{Q}_{o}$ is the mean observed discharge over $T$ days. NNSE values range from 0 (poor performance) to 1 (perfect fit), with a score of 0.5 indicating that the model’s predictions are equivalent to the time-averaged mean of the observations. In this study, the NNSE value is calculated for each gauge over time $T$, and the average NNSE across all gauges in the input data is reported.
$NNSE=\frac{\Sigma_{t=1}^{T}(Q_{m}^{t}-\overline{Q}_{o})^{2}}{\Sigma_{t=1}^{t}(Q_{o}^{t}-Q_{m}^{t})^{2}+(Q_{o}^{t}-\overline{Q}_{o})^{2}}\vspace{4mm}$ (11)
IV LSTM Benchmark Runs
IV-A Experiment Setup
All experiments are conducted using the model architecture shown in Fig. 3 and the activation functions listed in Table IV. The benchmark runs utilize Linear Space, Linear Time, and Annual Fourier Time encodings. Streamflow time series data is not trained but predicted as target.
IV-B Three-Nation Combined Benchmark Run
This run evaluates model performance on the combined CAMELS dataset from three nations: the US, UK, and Chile. The input data include static properties that are common across all three regions, whereas predicted targets consist of precipitation, mean temperature, and streamflow. Results are demonstrated in Table V.
TABLE V: CAMELS Three-Nations Combined Benchmark Run
@@ IV-B Three-Nation Combined Benchmark Run
IV-C CAMELS Caravan US Benchmark Runs
The runs compare model performance between the original CAMELS-US dataset and the US sub-dataset within the Caravan dataset. The CAMELS-US model is trained using selected static and time series data from the original CAMELS-US dataset, while the Caravan-US model is trained using selected static and time series data from the US sub-dataset within Caravan. For both runs, we utilize the same time series features—precipitation and mean temperature—for training and predict the same targets: precipitation, mean temperature, and streamflow. Results are presented in Table VI.
TABLE VI: CAMELS Caravan US Comparison
@@ IV-C CAMELS Caravan US Benchmark Runs
IV-D Caravan PCA Runs
This experiment explores the use of Principal Component Analysis (PCA) [48] to reduce the dimensionality of static properties in the Caravan input data. The Caravan sub-datasets contain over 200 static properties, nearly seven times more than the number of static input properties used in the original CAMELS studies. While this extensive range of static properties enhances model training, it significantly increases computational demands and GPU usage. To address this, we apply PCA, a widely adopted dimensionality reduction technique, to reduce the number of static input features to a level comparable with that in the CAMELS studies. In this experiment, we set the explained variance threshold to 90%, resulting in the reduction of static input features to approximately 30.
The first part of the experiment assesses the effect of PCA on model trained on US sub-dataset within the Caravan dataset, representing a small-scale input. The second part of the experiment examines the effect of PCA on model trained on North America regional data (HYSETS) [37] within the Caravan dataset, representing a large-scale input.
Results, shown in Table VII and Table VIII, indicate that the models trained with static properties obtained from PCA perform comparably to the models trained with original static properties, demonstrating that the reduction in input static dimensionality does not significantly compromise model accuracy. These findings validate PCA as an effective approach for train hydrology time series models with high static dimensionality.
TABLE VII: Caravan US PCA Experiment
@@ IV-D Caravan PCA Runs
TABLE VIII: Caravan Hysets PCA Experiment
@@
IV-E Caravan Global Benchmark Run
This study provides insights into the model’s applicability to large-scale global hydrology datasets. The input data is compiled by concatenating all seven Caravan sub-datasets, which encompass catchments from four continents. As shown in Table IX, accuracy decreases slightly compared to small-scale, individual nation fits (Table VI), yet overall model performance remains relatively high, demonstrating its robustness on global datasets.
TABLE IX: Caravan Global Fit
@@ IV-E Caravan Global Benchmark Run
V Static Properties and Spatial Temporal Encodings Experiment
@@
V-A Experiment Setup
This set of experiments examines the influence of static properties and spatial-temporal encodings on rainfall-runoff modeling accuracy, demonstrated through eight run configurations on four datasets. The first experiment uses the original CAMELS-US dataset [21], representing a traditional, locally-sourced small-scale dataset. Results are presented in Fig. 7. The second experiment utilizes the Three-Nation Combined CAMELS dataset [21, 22, 23], representing a locally-sourced mid-scale dataset. Results are shown in Fig. 7. The third experiment is conducted on the US sub-dataset of the Caravan dataset [36], representing a globally-sourced small-scale dataset. Results are shown in Fig. 7. The fourth experiment uses the full Caravan concatenated dataset [36], representing a globally-sourced large-scale dataset. Results are shown in Fig. 7. All four spatial-temporal encoding configurations tested are presented in Table X. For each configuration, we conduct runs with and without static features as input to highlight the impact of static features. The model architecture remained consistent across all runs.
TABLE X: Static Properties and Encodings Experiment Configurations
@@ V-B Experiment Findings
VI Comparison with Foundation Models
Recent studies in time series analysis have focused on a foundation model approach, inspired by the success of large language models [49]. Since 2022, there has been a spike in publications on pre-trained time series foundation models, with the majority built on Transformers [50] and multi-layer perceptrons (MLPs) [51] architectures. Some models support the use of static exogenous variables as input [52, 53], while others only accept time series input [54].
@@ VI Comparison with Foundation Models
In this experiment, we present a performance comparison between our LSTM-based rainfall-runoff model and the TSMixer foundation model [52] on CAMELS-US [21] data. We leverage the Nixtla Neuralforecast framework [55] to train the TSMixerx model, a variant of the TSMixer model that allows static exogenous features along with multivariate time series as input. The results, shown in Fig. 11, Fig. 11, Fig. 11, and Fig. 11, demonstrate that the LSTM model outperforms the TSMixerx model. In future work, we will expand this by looking at other time series models and a new Foundation model MultiFoundationPattern discussed for earthquakes[56].
VII Conclusion and Future Work

### v2

@@ 3.2 Model Setup
Table 4: Model Activation Setup
@@
3.3 Model Training and Evaluation
Inputs to the model can be classified into known inputs and observed inputs Fox et al. (2022). Known inputs refer to features that are known in both the past and future. In this hydrology study, static known inputs are the exogenous features such as climatic signatures, hydrologic signatures, and catchment topography. Additionally, time series known inputs exist in the form of seasonal patterns hidden in all time series features. Observed inputs are features known only for the past time periods but unknown in the future. In hydrology, those features include precipitation, temperature, and streamflow. Model output, known as targets, are the time-dependent predicted properties. In this study, predicted targets are precipitation, temperature, and streamflow.
@@ 3.3 Model Training and Evaluation
Spatial and Temporal Encodings. Like many fields of science, hydrology time series data exhibit strong seasonal patterns. It can reasonably be assumed that, at a specific gauge location, the precipitation and streamflow in October of one year will be similar to those recorded in October of the previous year. Furthermore, research has identified the approximate water residence times in various reservoirs, such as rivers, lakes, soil, and the atmosphere Pidwirny (2006). To effectively capture these known dependencies within the time series properties, spatial and temporal encodings are incorporated into the model during training.
@@
NNSE. Model fit is quantified using normalized Nash-Sutcliffe Efficiency (NNSE) scores Nossent and Bauwens (2012), calculated using Equation 11, where $T$ represents the total number of days, $Q_{m}^{t}$ represents the modeled discharge on day $t$, $Q_{o}^{t}$ represents the observed discharge on day $t$, and $\overline{Q}_{o}$ is the mean observed discharge over $T$ days. NNSE values range from 0 (poor performance) to 1 (perfect fit), with a score of 0.5 indicating that the model’s predictions are equivalent to the time-averaged mean of the observations. In this study, the NNSE value is calculated for each gauge over time $T$, and the average NNSE across all gauges in the input data is reported.
$NNSE=\frac{\Sigma_{t=1}^{T}(Q_{m}^{t}-\overline{Q}_{o})^{2}}{\Sigma_{t=1}^{t}(Q_{o}^{t}-Q_{m}^{t})^{2}+(Q_{o}^{t}-\overline{Q}_{o})^{2}}$ (11)
4 LSTM Benchmark Runs
4.1 Experiment Setup
All experiments are conducted using the model architecture shown in Fig. 3 and the activation functions listed in Table 4. The benchmark runs utilize Linear Space, Linear Time, and Annual Fourier Time encodings. We leverage multivariate time series forecasting RAMAN and SUNILKUMAR (1995). Streamflow time series data is not trained but predicted as target.
4.2 Three-Nation Combined Benchmark Run
This run evaluates model performance on the combined CAMELS dataset from three nations: the US, UK, and Chile. The input data include static properties that are common across all three regions, whereas predicted targets consist of precipitation, mean temperature, and streamflow. Results are demonstrated in Table 5.
Table 5: CAMELS Three-Nations Combined Benchmark Run
@@ 4.2 Three-Nation Combined Benchmark Run
4.3 CAMELS Caravan US Benchmark Runs
The runs compare model performance between the original CAMELS-US dataset and the US sub-dataset within the Caravan dataset. The CAMELS-US model is trained using selected static and time series data from the original CAMELS-US dataset, while the Caravan-US model is trained using selected static and time series data from the US sub-dataset within Caravan. For both runs, we utilize the same time series features—precipitation and mean temperature—for training and predict the same targets: precipitation, mean temperature, and streamflow. Results are presented in Table 6.
Table 6: CAMELS Caravan US Comparison
@@ 4.3 CAMELS Caravan US Benchmark Runs
4.4 Caravan PCA Runs
This experiment explores the use of Principal Component Analysis (PCA) Wold et al. (1987) to reduce the dimensionality of static properties in the Caravan input data. The Caravan sub-datasets contain over 200 static properties, nearly seven times more than the number of static input properties used in the original CAMELS studies. While this extensive range of static properties enhances model training, it significantly increases computational demands and GPU usage. To address this, we apply PCA, a widely adopted dimensionality reduction technique, to reduce the number of static input features to a level comparable with that in the CAMELS studies. In this experiment, we set the explained variance threshold to 90%, resulting in the reduction of static input features to approximately 30.
The first part of the experiment assesses the effect of PCA on model trained on US sub-dataset within the Caravan dataset, representing a small-scale input. The second part of the experiment examines the effect of PCA on model trained on North America regional data (HYSETS) Arsenault et al. (2020) within the Caravan dataset, representing a large-scale input.
Results, shown in Table 7 and Table 8, indicate that the models trained with static properties obtained from PCA perform comparably to the models trained with original static properties, demonstrating that the reduction in input static dimensionality does not significantly compromise model accuracy. These findings validate PCA as an effective approach for train hydrology time series models with high static dimensionality.
Table 7: Caravan US PCA Experiment
@@ 4.4 Caravan PCA Runs
Table 8: Caravan Hysets PCA Experiment
@@
4.5 Caravan Global Benchmark Run
This study provides insights into the model’s applicability to large-scale global hydrology datasets. The input data is compiled by concatenating all seven Caravan sub-datasets, which encompass catchments from four continents. As shown in Table 9, accuracy decreases slightly compared to small-scale, individual nation fits (Table 6), yet overall model performance remains relatively high, demonstrating its robustness on global datasets.
Table 9: Caravan Global Fit
@@ 4.5 Caravan Global Benchmark Run
5 Static Properties and Spatial Temporal Encodings Experiment
@@
5.1 Experiment Setup
This set of experiments examines the influence of static properties and spatial-temporal encodings on rainfall-runoff modeling accuracy, demonstrated through eight run configurations on four datasets. The first experiment uses the original CAMELS-US dataset Addor et al. (2017), representing a traditional, locally-sourced small-scale dataset. Results are presented in Fig. 7. The second experiment utilizes the Three-Nation Combined CAMELS dataset Addor et al. (2017); Coxon et al. (2020); Alvarez-Garreton et al. (2018), representing a locally-sourced mid-scale dataset. Results are shown in Fig. 7. The third experiment is conducted on the US sub-dataset of the Caravan dataset Kratzert et al. (2023), representing a globally-sourced small-scale dataset. Results are shown in Fig. 7. The fourth experiment uses the full Caravan concatenated dataset Kratzert et al. (2023), representing a globally-sourced large-scale dataset. Results are shown in Fig. 7. All four spatial-temporal encoding configurations tested are presented in Table 10. For each configuration, we conduct runs with and without static features as input to highlight the impact of static features. The model architecture remained consistent across all runs.
Table 10: Static Properties and Encodings Experiment Configurations
@@ 5.2 Experiment Findings
6 Data Splitting and Cube Root Normalization Experiment
Table 11 summarizes four experimental configurations for training and evaluating the LSTM model on precipitation, varying the train/validation set splitting (time-based vs. location-based) and preprocessing (just min-max vs. min-max and cube root normalization). The results, sorted by the Normalized Nash–Sutcliffe Efficiency (NNSE), shed light on how domain-specific considerations can significantly affect model performance.
Table 11: Performance of LSTM across different data preprocessing and splitting strategies.
Splitting Preprocessing MSE MAE NNSE
Time min-max 0.00035 0.00949 0.6177
Time min-max & cube root 0.00682 0.06076 0.7180
Location min-max 0.00020 0.00707 0.7443
Location min-max & cube root 0.00365 0.04121 0.8240
Location-based splitting with cube root normalization achieves the highest NNSE (0.8240), demonstrating that LSTMs perform better in modeling rainfall-runoff relationships for unseen catchments when precipitation data are cube-root-transformed. This transformation reduces skewness caused by large variations in precipitation across catchments with differing climates and land surfaces, enabling better learning from extreme values. In comparison, using raw data with the same location-based split results in a lower NNSE (0.7443), likely due to skewed values disrupting stable training across diverse catchments.
Time-based and location-based splitting highlight different objectives in hydrological modeling. Time-based splitting aims at forecasting future runoff within the same catchments, making it suitable for operational predictions, but yields lower NNSE values (0.6177–0.7180). Location-based splitting evaluates a model’s ability to generalize to new catchments, which is essential for global-scale or unmonitored applications. The higher NNSE for location-based splitting suggests it better captures robust spatial patterns.
MSE and MAE, reported in Table 11, vary across test sets (distinct years for time-based, distinct catchments for location-based) and are not directly comparable. NNSE, however, normalizes residuals by the observed variance, providing a consistent and reliable measure of model performance across different configurations. This makes NNSE the most appropriate metric for evaluating how well the model captures hydrological dynamics.
7 Comparison with other Pattern and Foundation Models
Recent advancements in time series analysis have adopted a foundation model approach, drawing inspiration from large language models Mirchandani et al. (2023). Since 2022, there has been a surge in research on foundation models for time series forecasting, primarily utilizing Transformer Vaswani et al. (2023) and MLP Tolstikhin et al. (2021) architectures. Some models incorporate static exogenous variables as input Chen et al. (2023); Das et al. (2023), while others rely solely on time series data Ansari et al. (2024). These models aim to generalize well in data-scarce tasks, integrate multimodal data, and offer explainable decision-making processes Ye et al. (2024). Research in this area focuses on two strategies: pre-training foundation models with time series data and adapting pre-trained language models for time series tasks Ye et al. (2024). Their success stems from effectively capturing multimodal distributions and repeated seasonal patterns Gruver et al. (2024).
This study compares the performance of state-of-the-art foundation models and our LSTM-based pattern model on the hydrology rainfall-runoff task. Pattern models, smaller-scale models trained for specific tasks Jafari et al. (2024), are contrasted with broad-purpose foundation models. All benchmarks use CAMELS-US Addor et al. (2017) data and the Nixtla Neuralforecast framework Olivares et al. (2022). Pattern models include TFT Lim et al. (2021), TCN Bai et al. (2018), DilatedRNN Chang et al. (2017), GNN-Coder Jafari et al. (2024), TimesNet Wu et al. (2022), and TiDE Das et al. (2023). Foundation models analyzed include TSMixer Chen et al. (2023), iTransformer Liu et al. (2023), and PatchTST Nie et al. (2022), pre-trained on datasets like TrafficL Lai et al. (2018), M4 Makridakis et al. (2020), and Weather Wu et al. (2021). Adapted models from LLMs, such as Chronos Ansari et al. (2024), are also evaluated, with Chronos using T5 Raffel et al. (2020) and GPT-2 Radford et al. (2019) backbones.
Results are presented through two experiments: one using min-max and cube root normalization and the other using raw precipitation data.
@@ 7 Comparison with other Pattern and Foundation Models
7.1 Model Performance Comparison of Processed Data Forecasting (Spatial Validation)
Our LSTM model is trained under benchmark configurations in both univariate and multivariate settings, using all 6 time series properties. Missing streamflow values are imputed with the mean of all observed streamflow data for both univariate and multivariate training. We apply min-max and cube root normalization on precipitation and streamflow data. All foundation and pattern models are trained and evaluated on the same training and validation locations, with a sequence length of 21 days and a forecast horizon of 1 day. Static exogenous features are utilized if supported by the model. As shown by the mean squared error results in Table 12, the multivariately trained LSTM model outperforms all other approaches under these benchmark conditions. Note that this study did not study the best hyperparameters that yield the best results for each model.
Model fits shown in Fig. 11, Fig. 11, Fig. 11, and Fig. 11 illustrate LSTM model performance compared to the TSMixer model Chen et al. (2023) that allows static exogenous features along with multivariate time series as input.
Table 12: Performance comparison of various models
Model Type a Input&Output b P c SR d Tmax e Tmin f VP g Q h
LSTM-multivariate P MV&MV 0.00298 0.00259 0.00019 0.00039 0.00015 0.00026
LSTM-univariate P UV&UV 0.00340 0.00321 0.00038 0.00038 0.00024 0.00036
TFT P UV&UV 0.00909 0.00843 0.00077 0.00093 0.00066 0.05686
TSMixer-M4 F MV&MV 0.00721 0.00599 0.00083 0.00051 0.00097 0.00042
TSMixer-TrafficL F MV&MV 0.00738 0.00609 0.00086 0.00053 0.00101 0.00044
TSMixer F MV&MV 0.00866 0.00698 0.00097 0.00073 0.00134 0.00065
TCN P UV&UV 0.01074 0.00889 0.00118 0.00108 0.00147 0.00286
iTransformer-M4 F MV&MV 0.00796 0.00745 0.00125 0.00080 0.00127 0.00041
iTransformer F MV&MV 0.00756 0.00776 0.00127 0.00066 0.00123 0.00054
DilatedRNN P UV&UV 0.00986 0.00896 0.00128 0.00098 0.00141 0.00078
iTransformer-TrafficL F MV&MV 0.00847 0.00790 0.00140 0.00094 0.00139 0.00042
GNNCoder-univariate P UV&UV 0.00710 0.00781 0.00151 0.00119 0.00156 0.00040
PatchTST-Weather F UV&UV 0.00892 0.00847 0.00169 0.00112 0.00164 0.00043
PatchTST-TrafficL F UV&UV 0.00890 0.00847 0.00169 0.00111 0.00165 0.00043
PatchTST-M4 F UV&UV 0.00892 0.00846 0.00169 0.00111 0.00166 0.00043
GNNCoder-multivariate P MV&MV 0.00861 0.00166 0.00171 0.00798 0.00247 0.00040
TimesNet P UV&UV 0.00939 0.00872 0.00176 0.00119 0.00177 0.08661
PatchTST F UV&UV 0.00952 0.00868 0.00184 0.00112 0.00167 0.00055
TiDE P UV&UV 0.01028 0.00870 0.00204 0.00160 0.00268 0.00096
Chronos-T5 F UV&UV 0.03590 0.01527 0.00628 0.00567 0.00724 0.01969
Chronos-GPT2 F UV&UV 0.02804 0.01261 0.00631 0.00663 0.02564 0.01423
aPattern or foundation model, bUniVariate or MultiVariate training and forecasting
cPrecipitation, dSolar radiation, eMaximum temperature, fMinimum temperature, gVapor pressure, hStreamflow.
Table shows mean squared error results, with the lowest for each predicted target highlighted in red.
Most foundation models prioritize univariate forecasting for broad applicability Ansari et al. (2024). However, a multivariate approach, which jointly trains and forecasts correlated time series properties, can improve performance. Using CAMELS-US data Addor et al. (2017), we found that multivariate training outperformed univariate methods, particularly in tasks like rainfall-runoff modeling where correlations between properties are crucial.
7.2 Model Performance Comparison of Raw Precipitation Forecasting (Temporal Validation)
The evaluation of raw precipitation data reveals notable performance differences among the tested models, as summarized in Table 13 (MSE, MAE, and NNSE).
The MultiFoundationCore achieved the best results (MSE: 26.8345, MAE: 2.4773, NNSE: 0.6312), closely followed by the MultiFoundationPattern (MSE: 26.9578, MAE: 2.5434, NNSE: 0.6305). These advanced architectures effectively model temporal patterns in precipitation, making them robust solutions for hydrological applications.
Among baseline models, LSTM and TFT delivered strong results, with LSTM achieving an NNSE of 0.6305 (MSE: 27.0633, MAE: 2.5089) and TFT slightly behind with an NNSE of 0.6294 (MSE: 27.1544, MAE: 2.3527). Other models, such as TimeLLM (NNSE: 0.6021) and TCN (NNSE: 0.6041), captured temporal dependencies effectively but lagged behind the more advanced architectures.
Foundation models like TSMixer-M4 (MSE: 35.8996) and iTransformer-M4 (MSE: 33.0192) struggled to adapt to raw precipitation data, underscoring the need for domain-specific adjustments. Hybrid models such as DilatedRNN performed competitively (MSE: 27.3794, NNSE: 0.6275), while models like PatchTST-M4 and TiDE showed moderate results, with NNSE values of 0.6096 and 0.6071, respectively, indicating limitations in capturing the complexities of raw precipitation data.
Table 13: Performance of time-segmented precipitation forecasting models, where “P” shows pattern models, “F” indicates foundation models, “MV” stands for multivariate inputs/outputs, and “UV” for univariate.
Model Type Input&Output MSE MAE NNSE
TSMixer-M4 F UV&UV 35.8996 3.1405 0.5623
iTransformer-M4 F UV&UV 33.0192 2.8941 0.5827
iTransformer-TrafficL F UV&UV 32.8197 2.7774 0.5842
TSMixer-TrafficL F UV&UV 31.9283 2.8479 0.5909
PatchTST-Weather F UV&UV 31.3607 2.7944 0.5952
TimesNet P MV&UV 30.7948 2.7750 0.5996
TimeLLM F UV&UV 30.4791 2.6698 0.6021
TCN P MV&UV 30.2239 2.6142 0.6041
PatchTST-TrafficL F UV&UV 29.8829 2.7299 0.6068
TiDE P MV&UV 29.8450 2.8262 0.6071
PatchTST-M4 F UV&UV 29.5362 2.7537 0.6096
DilatedRNN P MV&UV 27.3794 2.5771 0.6275
TFT P MV&UV 27.1544 2.3527 0.6294
LSTM P MV&UV 27.0633 2.5089 0.6302
MultiFoundationPattern F MV&UV 26.9578 2.5434 0.6305
MultiFoundationCore F MV&UV 26.8345 2.4773 0.6312
The MultiFoundationCore architecture demonstrates strong adaptability and scalability in hydrology, effectively capturing intricate patterns across various preprocessing techniques and data-splitting strategies. Its robust performance underscores its suitability for diverse hydrological applications
8 MultiFoundationCore Model
The MultiFoundationCore framework enhances time series forecasting in complex scientific domains by integrating multiple pre-trained foundation models and specialized pattern models. Building on the MultiFoundationPattern Jafari et al. (2024), it uses advanced cross-attention mechanisms to synthesize diverse inputs, capturing intricate spatial and temporal dependencies. This scalable approach leverages the strengths of various models to improve prediction accuracy.
Initial results in the hydrology domain show significant improvements in forecasting accuracy, particularly for complex spatiotemporal patterns. These findings underscore the framework’s potential to generalize across scientific applications, advancing reliable time series forecasting in challenging domains.
9 Conclusion and Future Work